Naledi leaves home in Durban at 9:00 a.m. and drives towards PMB on N3 at

80 km/h. Susan, in PMB 80 km from Durban, leaves at the same time and

travels towards Durban on N3 as 60 km/h. Where will they meet for lunch?


1
Expert's answer
2021-06-13T17:35:34-0400
"t_1=t_2\\\frac{x}{v_1}=\frac{d-x}{v_2}
\\\frac{x}{80}=\frac{80-x}{60}\\x=45.7\ km"